OpenStudy (anonymous):

What are the square roots of 256?

OpenStudy (anonymous):

it is 16.

OpenStudy (anonymous):

he's right ^^^ 256 = 2^8

OpenStudy (anonymous):

so 2^4 = 16

OpenStudy (anonymous):

16*16=256